OpenStudy (s):

Find the general solution of the differential equation if x is restricted to the interval (0, infinity). (x^2)y''-2xy'-4y=0

OpenStudy (anonymous):

That's called cauchy-euler (r^2-r)-2(r)-4=0 solve for r r=-1, 4

OpenStudy (anonymous):

y= C t^-1 + C t^4

OpenStudy (anonymous):

t should be x

OpenStudy (s):

Ok I think I get it thanks! The only thing, why is it (r^2-r)?

OpenStudy (anonymous):

we initially assume that the solution must be y= t^r y'= r t^(r-1) y''= r(r-1) t^(r-2)

OpenStudy (s):

Ooooo ok that makes sense. Thanks a lot!

OpenStudy (anonymous):

Substitute \(x=e^t \implies t=\ln{x}\), then: \[\frac{dy}{dx}={1 \over x}{dy \over dt} \text{ and } {d^2y \over dx^2}={1 \over x^2}({d^2y \over dt^2}-{dy \over dy})\] Plug this into your original DE you get: \[y''(t)-y'(t)-2y''(t)-4y(t)=0 \implies y''(t)-3y'(t)-4y(t)=0.\] The auxiliary equation of this HDE is \(m^2- 3m-4=(m-4)(m+1)=0\). Then the solution is \(y(t)=c_1e^{4t}+c_2e^{-t}\). Re-substitute with \(t=\ln{x}\) and that should give you the general solution: \[y(x)=c_1x^4+c_2x^{-1}.\]

OpenStudy (anonymous):

imrans solution seems easier.

OpenStudy (s):

I get this one as well. Thanks a lot for your help. Now I'm gonna refer to this and try to do similar problems. Thanks!